OpenStudy (anonymous):

Someone please help: A student found the exact value of cos 30 degrees using the fact that 30 degrees = 60 degrees minus 30 degrees and got zero. What was his error?

OpenStudy (anonymous):

\[\cos 30=\cos \left( 60-30 \right)=\cos 60\cos 30+\sin 60\sin 30\] the student used wrong formula and written - sign in place of + sign as given above.